OpenStudy (anonymous):

Solve the following system of equations. 4x + 7y – 2z = 0 3x – 5y + 3z = 9 3x + 6y – z = 1 Choices: A.(1, 0, 2) B.(–1, 0, –2) C. (1, 0, –2) D.(–1, 0, 2)

OpenStudy (anonymous):

do you want help or just an answer?

OpenStudy (anonymous):

can you do both please?

OpenStudy (anonymous):

Subtract the second equation from the third. That'll eliminate your x variable

OpenStudy (anonymous):

oh ok thank you

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

I'm still working it out

OpenStudy (anonymous):

is it C?

OpenStudy (anonymous):

Close, but no

OpenStudy (anonymous):

Lets assume your x and y are correct.

OpenStudy (anonymous):

If we plug into 3x + 6y – z = 1 We have: 3 + 0 - z = 1 3 - z = 1 -z = -2 z = 2

OpenStudy (anonymous):

But if we assume that z and y are correct: 3x + 0 - (-2) = 1 3x + 2 = 1 3x = -1 x = -1/3

OpenStudy (anonymous):

So there's something wrong here.

OpenStudy (anonymous):

Can you show some of your work and we can see where you went wrong?

OpenStudy (anonymous):

ok I subtracted the second and third equation to get -11y+5z=10

OpenStudy (anonymous):

then I got kind of confused from there

OpenStudy (anonymous):

Oh, well that's not quite right. 3x -5y + 3z = 9 - (3x + 6y -z) = -1 ============ 0x -11y +4z = 8

OpenStudy (anonymous):

oh I got it wrong from the beginning

OpenStudy (anonymous):

Actually though, as tempting as that is. I'm going to change my vote. Lets eliminate the z's

OpenStudy (anonymous):

That might be eaiser: Lets subtract 3 times the 3rd equation from the second one.

OpenStudy (anonymous):

is it D?

OpenStudy (anonymous):

A(1,0,2)